please help me translate the following:
Six less than three times a number is equal to the difference of fourteen and twice a number.
Thank You

Answer by macston(5194)

Let the number be N
six less than three times N is 3N-6
This equals the difference of fourteen and twice the number or 14-2N
So the equation is 3N-6=14-2N ANSWER This is the equation described above
3N-6=14-2N Add 2N to each side
5N-6=14 Add six to each side
5N=20 Divide each side by 5
N=4